How do I do this problem? Functions.
If f(x) = 2x^2 + 17,
find the following.
f(2)
f(sqrt 17)
f(2 + sqrt 17)
f(2) + f(sqrt 17)
f(4x)

Answer by waynest(281)

f(2) = 2(2)^2 + 17
= 2(4) + 17
= 8 + 17
= 25
f(sqrt17) = 2(sqrt17)^2 + 17
= 2(17) + 17
= 34 + 17
= 51
f(2 + sqrt17) = 2(2 + sqrt17)^2 + 17
= 2(4 + 4sqrt17 + 17) + 17
= 8 + 8sqrt17 + 34 + 17
= 59 + 8sqrt17
f(2) + f(sqrt) = (2(2)^2 + 17) + (2(sqrt17)^2 + 17)
= 25 + 51
= 76
f(4x) = 2(4x)^2 + 17
= 2(16x^2) + 17
= 32x^2 + 17